General Physics

a body has a weight W kg on the surface of earth.what will be its weight on the surface of moon in gravitational units of force and si units

vinayak khandelwal
13 Years agoGrade

4 Answers

tejas naresh patel
13 Years ago

mgmoon

Chetan Mandayam Nayakar
13 Years ago

acceleration due to gravity on surface of moon=g/6

answer=Wg/6 N

Akash Kumar Dutta
13 Years ago

mass of body=W/g
at moon weight= mg/6=W/6
g-Force= W/6g    (g-Force)ANS

tejas naresh patel
13 Years ago

MG/6